# A Survey on Recommendation Unlearning: Fundamentals, Taxonomy, Evaluation, and Open Questions

### v1

@@ IV Evaluation
Based on the design principles outlined above, the evaluation of recommendation unlearning methods primarily focuses on three key aspects: unlearning completeness, unlearning efficiency, and model utility. Each of these aspects plays a crucial role in evaluating the effectiveness and practicality of unlearning techniques in real-world recommender systems. In this section, we provide a comprehensive overview of the evaluation process, detailing the commonly used datasets, recommendation models, and evaluation metrics employed to assess these three aspects.
@@ IV-A Datasets
  • Yelp [^2] https://www.yelp.com/dataset: The Yelp dataset was originally compiled for the Yelp Dataset Challenge and contains users’ reviews of restaurants [5]. The company Yelp [^3] https://www.yelp.com is a platform that publishes crowd-sourced reviews of restaurants. which is a chance for students to conduct research or analysis on Yelp’s data and share their discoveries.
@@
  • BookCrossing [^6] http://www.informatik.uni-freiburg.de/~cziegler/BX: The BookCrossing dataset was collected from the Book-Crossing community and contains book ratings.
@@
  • MIND [110]: The MIND dataset contains user click logs from Microsoft News, along with associated textual news information.
@@
  • LFM-2B [^9] https://www.cp.jku.at/datasets/LFM-2b: The LFM-2B dataset contains over 2 billion listening events, designed for music retrieval and recommendation tasks [77, 88]. It also includes user attributes such as gender, age, and country.
@@
IV-B Models
To verify the effectiveness of the proposed methods, recommendation unlearning techniques are often evaluated on various collaborative filtering models. In addition to the specific models or scenarios mentioned in Section III, we list the widely used collaborative filtering model structures as follows:
  • MF [48]: The traditional matrix factorization method.
  • BPR [84]: A classic recommendation model that optimizes matrix factorization using the Bayesian Personalized Ranking (BPR) objective function.
  • WMF [38, 17]: A non-sampling recommendation model that treats all missing interactions as negative instances, applying uniform weighting to them.
  • DMF [116]: A representative deep learning-based recommendation model that builds on matrix factorization.
  • NCF [34]: A key collaborative filtering model that leverages neural network architectures.
  • MultVAE [60]: A collaborative filtering model that learns embeddings by decoding the variational encoding.
  • LightGCN [33]: A state-of-the-art collaborative filtering model that simplifies graph convolution networks to enhance recommendation performance.
IV-C Metrics
IV-C1 Unlearning performance
@@
As mentioned in Section II-D, there are four perspectives of completeness. We introduce the metrics from each perspective as follows:
  • Algorithmic perspective: From this perspective, complete unlearning is defined solely as retraining the model from scratch, making the evaluation of algorithmic completeness largely theoretical. Exact unlearning naturally achieves full completeness, while other methods do not.
  • Parametric perspective: Although there is no literature that directly evaluates the parametric similarity between the unlearned and retrained models, some research leverages Membership Inference Attacks (MIA) to assess completeness. In this context, the input to the attacker is typically embeddings (i.e., parameters) from the recommendation model. Thus, we consider MIA evaluation as a measure of parametric completeness. Specifically, MIA aims to infer whether a data record was part of the model’s training set. If MIA determines that the unlearned data is no longer in the training set, unlearning is considered complete. Therefore, MIA evaluation uses the attacker’s performance as a metric, with typical evaluation metrics including accuracy, recall, precision, Area Under the Curve (AUC), and F1 score.
  • Functional perspective: This perspective requires the model to behave or perform similarly to the retrained model. Therefore, the evaluation of completeness from a functional perspective is based on recommendation performance. Representative metrics commonly used include Normalized Discounted Cumulative Gain (NDCG), Hit Ratio (HR) [32], and loss values such as Root Mean Square Error (RMSE). The evaluation is typically twofold: one on the forgot set and one on the retain set. Specifically, a complete unlearning from the functional perspective should result in low performance on the forgotten set, while maintaining high performance on the retained set.
  • Attack perspective: Backdoor attacks are a widely used method for evaluating completeness in traditional machine unlearning [69, 72]. In the context of recommendation tasks, where their data consists of user-item interactions (e.g., explicit feedback with rating values or implicit feedback with binary indicators), injecting a backdoor at the sample level is challenging. To address this, some research conducts poisoning attacks by flipping labels to create negative data. If unlearning is completely achieved, the negative impact of the flipped data is removed, leading to an improvement in recommendation performance, which aligns with the functional perspective of evaluation. In the case of attribute unlearning, the primary goal is to make the attribute distinguishable to attackers. Therefore, completeness evaluation of attribute unlearning is typically conducted from an attack perspective, using classification metrics mentioned above, e.g., AUC and F1 score.
@@
IV-C2 Other Deeper Impact
@@
Li et al. [16] highlight an important concern: exact unlearning, when applied to recommender systems, can inadvertently exacerbate fairness issues. Exact unlearning typically relies on an ensemble retraining framework, where similar users (or items) are grouped together within the same data subset for retraining. This process leads to a phenomenon where advanced users, i.e., those with richer interaction histories, tend to be clustered together, while disadvantaged users, who may have less frequent interactions, are also grouped separately. As a result, exact unlearning can unintentionally increase the disparity between these two groups, widening the gap in fairness. The outcome is that the system may disproportionately benefit advanced users, while disadvantaging those with fewer data records or lower engagement.
@@ V Discussion and Future Directions
Based on the presented taxonomy, we provide a comprehensive summary of the similarities and differences among various recommendation unlearning methods (i.e., exact unlearning vs. approximate unlearning), highlighting the key trends and developments in existing approaches. This summary inspires further research to explore two key questions in recommendation unlearning and, more broadly, in machine unlearning: What to unlearn? and how to unlearn? We also compare recommendation unlearning methods with general machine unlearning techniques, examining the similarities in the design principles as well as the distinct challenges posed by the recommendation domain. This comparison helps to clarify the specific requirements and constraints that recommender systems face, which may not be as prevalent in other machine learning tasks. Finally, we identify several open research questions in the field.
@@ V-A2 Unlearning Methods (how to unlearn)
In the context of unlearning in recommender systems, both exact and approximate unlearning approaches have been explored, with approximate methods gaining increasing popularity. This trend aligns with the ongoing developments in traditional machine unlearning research. The growing preference for approximate unlearning can be attributed to several factors, primarily the trade-off between theoretical guarantees and empirical performance. While exact unlearning provides a strong theoretical foundation for ensuring complete unlearning of forget set, its empirical performance often lags behind that of approximate methods [16]. This is particularly evident in real-world applications, where users typically prioritize practical performance (e.g., model utility and efficiency) over theoretical guarantees.
@@
V-A3 Rising Research Tasks
@@
These advancements highlight the growing need for unlearning methodologies that are not only model-specific but also adaptable to a broader range of use cases. As recommender systems become more sophisticated and pervasive, investigating recommendation unlearning across a variety of settings will be essential. The ongoing evolution of unlearning techniques promises to offer more robust, scalable solutions for handling sensitive and dynamic user data, while maintaining the functionality and fairness of recommender systems in increasingly complex environments.
@@ V-B1 Other Recommendation Scenarios
  • User-perspective unlearning. Most existing unlearning research adopts a system-centric perspective, where the recommender system itself is responsible for performing the unlearning operation. However, there is an emerging need to explore user-perspective unlearning, where users can proactively manage their data by requesting the deletion or modification of their preferences, without necessarily being dependent on the system to do so. This would enable users to erase their personal data or adjust their preferences autonomously, even in systems where the platform does not explicitly support such features. User-perspective unlearning has significant implications for user privacy and control, as it allows users to take ownership of their data and make unlearning requests directly, without waiting for the system to initiate or support the process. This scenario raises new challenges around how to empower users to manage their own data effectively, especially in platforms that are not inherently designed for such dynamic interactions.
@@ V-B2 Unlearning Audition
One of the central challenges in unlearning research is developing a standard, transferable method for evaluating completeness. This is particularly important in the context of sample-wise unlearning (i.e., unlearning a single piece of user-item interaction record). A key question is whether sample-wise unlearning is a genuine task. The memory or contribution of a single data point can be difficult to assess, as removing one or a few data points might not substantially affect the model’s accuracy, loss, or parameters due to the influence of other data points. Thus, developing metrics to accurately quantify the influence of sample-wise unlearning is critical. Alternatively, it may be impossible to assess sample-wise influence, with sample-wise unlearning existing only at an algorithmic level. This question is also present in traditional machine unlearning research.
@@ V-B3 Interpretability of Unlearning
  • What does a recommendation model learn, and what should be unlearned? Given the stochastic nature and randomness inherent in deep learning, this remains an open question. Prior studies provide definitions from an algorithmic perspective [42, 100, 109], but a more comprehensive definition that incorporates both the parametric and functional levels is desired. Such an approach would contribute significantly to the interpretability of both the learning and unlearning processes.
@@ V-B4 Attack and Defense
As unlearning techniques continue to evolve, so too do the potential vulnerabilities they may introduce [52]. While unlearning is designed to eliminate the influence of specific data points, it is not immune to adversarial threats. Robust unlearning must, therefore, incorporate strategies to safeguard the model from attacks that aim to exploit weaknesses in the unlearning process itself.
Moreover, unlearning can inadvertently create a false sense of security. In fact, unlearning the data that are perceived to be harmful or unnecessary may exaggerate the model’s vulnerability to future attacks. For example, the memorization of deep learning models exhibits an onion effect, where removing “outliers” from the training data may inadvertently increase the vulnerability of the “inner data” to attacks [13]. Furthermore, attackers can exploit the differences between the model parameters before and after unlearning to design more sophisticated and effective attacks [19].

### v2

@@ IV Evaluation
Based on the design principles outlined above, the evaluation of recommendation unlearning methods primarily focuses on three key aspects: unlearning completeness, unlearning efficiency, and model utility. Each of these aspects plays a crucial role in evaluating the effectiveness and practicality of unlearning techniques in real-world recommender systems. In this section, we provide a comprehensive overview of the evaluation process, detailing the commonly used datasets and evaluation metrics employed to assess these three aspects.
@@ IV-A Datasets
  • Yelp [^2] https://www.yelp.com/dataset: The Yelp dataset was originally compiled for the Yelp Dataset Challenge and contains users’ reviews of restaurants [asghar2016yelp]. The company Yelp [^3] https://www.yelp.com is a platform that publishes crowd-sourced reviews of restaurants. which is a chance for students to conduct research or analysis on Yelp’s data and share their discoveries.
@@
  • BookCrossing [^6] http://www.informatik.uni-freiburg.de/˜cziegler/BX: The BookCrossing dataset was collected from the Book-Crossing community and contains book ratings.
@@
  • MIND [wu2020mind]: The MIND dataset contains user click logs from Microsoft News, along with associated textual news information.
@@
  • LFM-2B [^9] https://www.cp.jku.at/datasets/LFM-2b: The LFM-2B dataset contains over 2 billion listening events, designed for music retrieval and recommendation tasks [melchiorre2021investigating, schedl2022lfm]. It also includes user attributes such as gender, age, and country.
@@
IV-B Metrics
IV-B1 Unlearning performance
@@
As mentioned in Section II-B, there are four perspectives of completeness. We introduce the metrics from each perspective as follows:
  • Algorithmic perspective: From this perspective, complete unlearning is defined solely as retraining the model from scratch, making the evaluation of algorithmic completeness largely theoretical. As mentioned in Section II-B, this definition can only be self-evaluated by providing training checkpoints. Thus, exact unlearning (i.e., retraining-based unlearning methods) naturally achieves algorithmic completeness, while other methods do not.
  • Parametric perspective: Although there is no literature that directly evaluates the parametric similarity between the unlearned and retrained models in recommendation unlearning, some research leverages Membership Inference Attacks (MIA) to assess completeness. In the context of recommendation unlearning, the input to the attacker is typically embeddings (i.e., parameters) from the recommendation model. Thus, we consider MIA evaluation as a measure of parametric completeness. Specifically, MIA aims to infer whether a data record was part of the model’s training set. If MIA determines that the unlearned data is no longer in the training set, unlearning is considered complete. Therefore, MIA evaluation uses the attacker’s performance as a metric, with typical evaluation metrics including accuracy, recall, precision, Area Under the Curve (AUC), and F1 score.
  • Functional perspective: This perspective requires the model to perform similarly to the retrained model. Therefore, the evaluation of completeness from a functional perspective is based on recommendation performance. Representative recommendation metrics commonly used include Normalized Discounted Cumulative Gain (NDCG), Hit Ratio (HR) [he2015trirank], and loss values such as Root Mean Square Error (RMSE). The evaluation is typically twofold: one on the forget set and one on the retain set. Specifically, a complete unlearning from the functional perspective should result in low performance on the forget set, while maintaining high performance on the retain set.
  • Attack perspective: Backdoor attacks are a widely used method for evaluating completeness in traditional machine unlearning [liu2022backdoor, liu2024backdoor]. In the context of recommendation tasks, where their data consists of user-item interactions (e.g., explicit feedback with rating values or implicit feedback with binary indicators), injecting a backdoor at the sample level is challenging. To address this, some research conducts poisoning attacks by flipping labels to create negative samples. If unlearning is completely achieved, the negative impact of these flipped samples is removed, leading to an improvement in recommendation performance, which aligns with the functional perspective of evaluation. In the case of attribute unlearning, the primary goal is to make the attribute distinguishable to attackers. Therefore, completeness evaluation of attribute unlearning is typically conducted from an attack perspective, using classification metrics mentioned above, e.g., AUC and F1 score.
@@
IV-B2 Other Deeper Impact
@@
Li et al. [CURE4Rec2024] highlight an important concern: exact unlearning, when applied to recommender systems, can inadvertently exacerbate fairness issues. Exact unlearning typically relies on an ensemble retraining framework, where similar users (or items) are grouped together within the same data subset for retraining. This process leads to a phenomenon where advanced users, i.e., those with richer interaction histories, tend to be clustered together, while disadvantaged users, who may have less frequent interactions, are also grouped separately. As a result, exact unlearning can unintentionally increase the disparity between these two groups, widening the gap in fairness. The outcome is that the system may disproportionately benefit advanced users, while disadvantaging those with fewer data records or lower engagement.
@@ V Discussion and Future Directions
Based on the presented taxonomy, we provide a comprehensive summary of the similarities and differences among various recommendation unlearning methods (i.e., exact unlearning vs. approximate unlearning), highlighting the key trends and developments in existing approaches. This summary inspires further research to explore two key questions in recommendation unlearning and, more broadly, in machine unlearning: what to unlearn? and how to unlearn? We also compare recommendation unlearning methods with general machine unlearning techniques, examining the similarities in the design principles as well as the distinct challenges posed by the recommendation domain. Finally, we identify several open research questions in the field, summarizing them into four aspects, i.e., other recommendation scenarios, unlearning audition, interpretability of unlearning, and attack and defense.
@@ V-A2 Unlearning Methods (how to unlearn)
In the context of unlearning in recommender systems, both exact and approximate unlearning approaches have been explored, with approximate methods gaining increasing popularity. This trend aligns with the ongoing developments in traditional machine unlearning research. The growing preference for approximate unlearning can be attributed to several factors, primarily the trade-off between theoretical guarantees and empirical performance. While exact unlearning provides a strong theoretical foundation for ensuring complete unlearning of forget set, its empirical performance often lags behind that of approximate methods [CURE4Rec2024]. This is particularly evident in real-world applications, where users typically prioritize practical performance (e.g., model utility and efficiency) over theoretical guarantees.
@@
V-A3 Emerging Research Tasks
@@
In summary, these advancements highlight the growing need for unlearning methodologies that are not only model-specific but also adaptable to a broader range of use cases. As recommender systems become more sophisticated and pervasive, investigating recommendation unlearning across a variety of settings will be essential. The ongoing evolution of unlearning techniques promises to offer more robust, scalable solutions for handling sensitive and dynamic user data, while maintaining the functionality and fairness of recommender systems in increasingly complex environments.
@@ V-B1 Other Recommendation Scenarios
  • User-perspective unlearning. Most existing unlearning research adopts a system-centric perspective, where the recommender system itself is responsible for performing the unlearning operation. However, there is an emerging need to explore user-perspective unlearning, where users can proactively manage their data by requesting the deletion or modification of their preferences, without necessarily being dependent on the system to do so. Several studies in the human-computer interaction community have explored similar concepts of user-perspective actions [liu2025filtering, shu2024rah]. This would enable users to erase their personal data or adjust their preferences autonomously, even in systems where the platform does not explicitly support such features. User-perspective unlearning has significant implications for user privacy and control, as it allows users to take ownership of their data and make unlearning requests directly, without waiting for the system to initiate or support the process. This scenario raises new challenges around how to empower users to manage their own data effectively, especially in platforms that are not inherently designed for such dynamic interactions.
@@ V-B2 Unlearning Audition
One of the central challenges in unlearning research is developing a standard, transferable method for evaluating completeness. This is particularly important in the context of sample-wise unlearning (i.e., unlearning a single piece of user-item interaction record). However, the memory or contribution of a single data point can be difficult to assess On the one hand, removing one or a few data points might not substantially affect the model’s accuracy, loss, or parameters due to the re-remember phenomenon illustrated in Figure 3. On the other hand, a small number of vital interactions can drastically impact particular embedding and recommendations due to the influence of other data points. Thus, developing metrics to accurately quantify the influence of sample-wise unlearning is critical. Alternatively, it may be impossible to assess sample-wise influence, with sample-wise unlearning existing only at an algorithmic level. This question is also present in traditional machine unlearning research.
@@ V-B3 Interpretability of Unlearning
  • What does a recommendation model learn, and what should be unlearned? Given the stochastic nature and randomness inherent in deep learning, this remains an open question. Prior studies provide definitions from an algorithmic perspective [jia2021proof, thudi2022necessity, weng2024proof], but a more comprehensive definition that incorporates both the parametric and functional levels is desired. Such an approach would contribute significantly to the interpretability of both the learning and unlearning processes.
@@ V-B4 Attack and Defense
As unlearning techniques continue to evolve, so too do the potential vulnerabilities they may introduce [leysen2023exploring]. While unlearning is designed to eliminate the influence of specific data points, it is not immune to adversarial threats. Robust unlearning must, therefore, incorporate strategies to safeguard the model from attacks that aim to exploit weaknesses in the unlearning process itself.
Moreover, unlearning can inadvertently create a false sense of security. In fact, unlearning the data that are perceived to be harmful or unnecessary may exaggerate the model’s vulnerability to future attacks. For example, the memorization of deep learning models exhibits an onion effect, where removing “outliers” from the training data may inadvertently increase the vulnerability of the “inner data” to attacks [carlini2022privacy]. Furthermore, attackers can exploit the differences between the model parameters before and after unlearning to design more sophisticated and effective attacks [chen2021machine].
@@ VI Conclusion
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (No. 62402148).
[Uncaptioned image] Yuyuan Li obtained his PhD degree in computer science from Zhejiang University, China, in 2023. He is currently an Associate Professor at Hangzhou Dianzi University. His research interests mainly focus on trustworthy machine learning. He has published over 30 papers in peer-reviewed journals and conferences, including Cell Patterns, NeurIPS, ICML, and ICLR.
[Uncaptioned image] Xiaohua Feng is currently pursuing a Ph.D. degree at the College of Computer Science and Technology, Zhejiang University, China. His undergraduate degree is from Wuhan University of Technology 2022. His research interests mainly focus on trustworthy machine learning. He has published papers in peer reviewed conferences and journals, including ICLR, WWW, EMNLP, and Cell Patterns.
[Uncaptioned image] Chaochao Chen (Senior Member, IEEE) obtained his PhD degree in computer science from Zhejiang University, China, in 2016, and he was a visiting scholar in University of Illinois at Urbana-Champaign, during 2014-2015. He is currently a Distinguished Research Fellow at Zhejiang University. Before that, he was a Staff Algorithm Engineer at Ant Group. His research mainly focuses on recommender system, privacy preserving machine learning, and graph machine learning. He has published more than 100 papers in peer reviewed journals and conferences.
[Uncaptioned image] Qiang Yang (Fellow, IEEE) received the graduation degree from Peking University and PhD degree from the Department of Computer Science, the University of Maryland, College Park, in 1989 and had been a faculty member, the University of Waterloo between 1989 and 1995. He is a fellow of Canadian Academy of Engineering (CAE) and Royal Society of Canada (RSC), former Chief Artificial Intelligence Officer of WeBank, a former chair professor of Computer Science and Engineering Department at Hong Kong University of Science and Technology (HKUST). He is currently a Chair Professor and Director of Hong Kong Polytechnic University, Academy for AI and a Chair Professor at Hong Kong University of Science and Technology (Guangzhou). He is the Conference Chair of AAAI-21, the Honorary vice president of Chinese Association for Artificial Intelligence(CAAI), the president of Hong Kong Society of Artificial Intelligence and Robotics(HKSAIR), the President of Investment Technology League (ITL) and Open Islands Privacy-Computing Open source Community. He is a fellow of AAAI, ACM, CAAI, IAPR, and AAAS. He was the Founding Editor in Chief of the ACM Transactions on Intelligent Systems and Technology (ACM TIST) and the Founding Editor in Chief of IEEE Transactions on Big Data (IEEE TBD). He received the ACM SIGKDD Distinguished Service Award, in 2017 and the Wu Wenjun outstanding contribution award of artificial intelligence, in 2019. His research interests are artificial intelligence, machine learning, data mining, and planning. He had been the Founding Director of the Huawei’s Noah’s Ark Research Lab between 2012 and 2015, the Founding Director of HKUST’s Big Data Institute, the Founder of 4Paradigm and the President of IJCAI (2017-2019). His latest books are Transfer Learning, Federated Learning, Privacy-preserving Computing, and Practicing Federated Learning.